There are 60% acid solution and 80% acid solution. Find the volume of each solution if both of them are mixed and the result is 20 liters of 65% acid solution.
Answer by richwmiller(17219)

By substitution
We know the total amount is 20 liters
We know the percentage of the total amount is 65%
We know the total amount has 13 liters of 100%
We want to know the amount of solution at 60%
We want to know the amount of solution at 80%
x+y=20,
0.6*x+0.8*y=0.65*20
0.6*x+0.8*y=13
x=20-y
0.6*(20-y)+0.8*y=13
12-0.6y+0.8*y=13
0.2*y=1
We now know amount of solution at 80%
y=5 liters at 80%
x=20-y
We now know amount of solution at 60%
x=15 liters at 60%
check
0.6*15+0.8*5=0.65*20
9+4=13
13=13
ok